OpenStudy (xoxolove456):

Jake and Kate can clean a room together in 5 hours. Working alone, jake can clean the room in 1 hour less time than Kate can. How long does it take Kate to clean the room alone?????

OpenStudy (xoxolove456):

Please help

OpenStudy (oswaldmurphy):

You got 5 hours. It takes Jake 1 less hour. So 5 = x + x - 1. 5 = 2x - 1 6 = 2x 3 = x 3 Is your answer :D